OpenStudy (raffle_snaffle):

For his trip, Davis deposited $24943 into an account paying 4% per year, compounded annually. Unfortunately, he had a medical emergency at the end of year 5 and had to withdraw $2389. At the end of year 7 he received a small inheritance and deposited $2116 into his account. How much did Davis have in his account at the end of 13 years?

OpenStudy (raffle_snaffle):

F = P(1+i)^n - 2389 + 2116

OpenStudy (raffle_snaffle):

@R.Hamza17

OpenStudy (raffle_snaffle):

@R.Hamza17

OpenStudy (raffle_snaffle):

The 4"% only pertains to the 24943?

OpenStudy (r.hamza17):

First of all you should find out the maturity value for 5 years. Than you should subtract the amount he had withdrawn from his account. Now on the new value again find the maturity amount for 2 more years. Now again add the amount he have got from his inheritance in the total amount. Noe after adding the amount again find the maturity amount for 6 more years than you will get the correct answer

OpenStudy (raffle_snaffle):

Okay so finding the maturity level I need to use F = P(1+i)^n?

OpenStudy (r.hamza17):

Step 1: P=24943 i=4% t=5

OpenStudy (r.hamza17):

Yes you have to use this formula

OpenStudy (raffle_snaffle):

okay I got 30346.97

OpenStudy (r.hamza17):

Step 2: A-2389=P Step 3: Using above P and the n=2 find A. Step 4: Now add 2116 into A=P Step 5: Now using this P and n=6 find A which will be your answer

OpenStudy (raffle_snaffle):

yes I get p = 30346-2389 = 27957.97

OpenStudy (raffle_snaffle):

Hey you can go help other ppl I am still working on it.

OpenStudy (r.hamza17):

Now follow the next steps to your answer

OpenStudy (raffle_snaffle):

What do you mean by using the above p?

OpenStudy (raffle_snaffle):

nvm got it

OpenStudy (raffle_snaffle):

I don't really quite understand how to solve step 2 because I don't know what I can do with A-2389=P

OpenStudy (raffle_snaffle):

I use F instead of A

OpenStudy (raffle_snaffle):

F-2389 = P SO the previous F I found is 30346.97334 then I solve for P?

OpenStudy (r.hamza17):

The value of P you got in the upper step use it and subtract the amount she have withdrawn.

OpenStudy (raffle_snaffle):

Step 1 I didn't solve for P I was given P

OpenStudy (raffle_snaffle):

F = p(1+i)^n A=p(1+i)^n

OpenStudy (r.hamza17):

In step 1 you were given P so you have to find A or F

OpenStudy (raffle_snaffle):

Okay which I did F = 30346.97

OpenStudy (r.hamza17):

Great now go for step 2

OpenStudy (raffle_snaffle):

30346 - 2389 = P

OpenStudy (r.hamza17):

Okay now after subtraction again use the formula to find it for the next 2 years

OpenStudy (raffle_snaffle):

at the same inerest, 4%

OpenStudy (raffle_snaffle):

I will BRB got to go to my instructors office hours for another class. I think I've got this one figure out. IF I have any further questions I will let you know.

OpenStudy (r.hamza17):

Okay. The interest rate will remain the same for all the steps which is 4%

OpenStudy (raffle_snaffle):

Hey I am back sorry for the late delay. I got caught up in other things after meeting up with my instructor.

OpenStudy (r.hamza17):

Okay so what can I help you with now?